OpenStudy (ffasinger):

If the area of the trapezoid below is 98 square units, what is the value of x?

OpenStudy (ffasinger):

@mathstudent55 please help

OpenStudy (decarr432):

Do you know how to find the area of a trapezoid?

OpenStudy (decarr432):

@FFASinger I am trying to help please stay on the question because currently it is saying you are just looking around..

OpenStudy (ffasinger):

sorry my computer lost intenet connection

OpenStudy (ffasinger):

im back @Decarr432

OpenStudy (decarr432):

Do you know how to find the area *Formula*?

OpenStudy (ffasinger):

i am gonna be honest no

OpenStudy (decarr432):

\[A= \frac{ B1+B2 }{ 2 }H\]

OpenStudy (ffasinger):

k

OpenStudy (decarr432):

we are finding H=X so what is B1+B2

OpenStudy (ffasinger):

24

OpenStudy (decarr432):

now 24/2?

OpenStudy (ffasinger):

12

OpenStudy (decarr432):

okay now 12*x=98 think of numbers you think will work for x to make it 98

OpenStudy (decarr432):

are there any answer choices given?

OpenStudy (ffasinger):

si hold on

OpenStudy (ffasinger):

a 14 b3.5 c 1.75d7

OpenStudy (decarr432):

okay what do you think will work?

OpenStudy (ffasinger):

14

OpenStudy (decarr432):

okay plug 14 in for X and solve what do you get?

OpenStudy (ffasinger):

ummmmm idk

OpenStudy (decarr432):

Please multiply 12*14 I am not doing this for you, you have to put forth the effort :)

OpenStudy (ffasinger):

ok well i did not know how to plug in the number so sorry . anyway i got 168

OpenStudy (decarr432):

Plugging in a number is like this x+2=3 When x=1 so you would do this (1)+2=3

OpenStudy (ffasinger):

cool thanks

OpenStudy (ffasinger):

k so i got 168

OpenStudy (ffasinger):

so thats not the ansewr

OpenStudy (decarr432):

Now is that the answer we need *hint* we need 98 :)

OpenStudy (ffasinger):

no

OpenStudy (decarr432):

So what number do you want to try next?

OpenStudy (ffasinger):

i tried them all and they did not work

OpenStudy (decarr432):

I don't see why not the answer choice I choose worked fine?

OpenStudy (decarr432):

okay lets try this then

OpenStudy (decarr432):

\[\frac{ 28*x }{ 2 }\]

OpenStudy (decarr432):

plug in x multiply then divide

OpenStudy (ffasinger):

where did you get 28

OpenStudy (decarr432):

Simply multiply x by 28 and divide by 2 :) try it with all of them

OpenStudy (decarr432):

@FFASinger